# A Survey of Robotic Language Grounding: Tradeoffs Between Symbols and Embeddings

### v1

@@ Abstract
With large language models, robots can understand language more flexibly and more capable than ever before. This survey reviews recent literature and situates it into a spectrum with two poles: 1) mapping between language and some manually defined formal representation of meaning, and 2) mapping between language and high-dimensional vector spaces that translate directly to low-level robot policy. Using a formal representation allows the meaning of the language to be precisely represented, limits the size of the learning problem, and leads to a framework for interpretability and formal safety guarantees. Methods that embed language and perceptual data into high-dimensional spaces avoid this manually specified symbolic structure and thus have the potential to be more general when fed enough data but require more data and computing to train. We discuss the benefits and tradeoffs of each approach and finish by providing directions for future work that achieves the best of both worlds.
Figure 1: Approaches to representing natural language for robotics fall along a spectrum from more symbol-like representations to more continuous embedding-like representations. However, most approaches use a mixture of both. SayCan (Ichter et al., 2022) uses a fixed ontology of predefined skills but implements these as neural value functions conditioned on language.
@@ 1 Introduction
Large language models (LLMs) have fueled a surge of interest in the problem of making robots understand natural language commands. Solving this problem requires mapping between words in language and actions or behaviors taken by the robot. Harnad (1990, 2007) defined the symbol grounding problem as constructing a mapping from symbols of a symbolic system or words in an utterance to sensorimotor substrates in the physical world. Large language models (LLMs), such as GPT-4 (OpenAI et al., 2023), semantically represent concepts without explicit higher-order symbols beyond the words in the text, leading many to try end-to-end approaches for robotic language understanding. Yet many recent works in robotic language understanding leverage large language models hand-in-hand with formal symbolic representations. For example, Code as Policies (Liang et al., 2023) generates Python code with predefined Python APIs. SayCan (Ichter et al., 2022) grounds natural language commands to predefined discrete skills implemented using a deep neural network. Other approaches take a more end-to-end approach, such as VIMA (Jiang et al., 2023), which learns a mapping from vision and language instructions to low-level robot actions such as joint states.
@@
This survey paper evaluates work that grounds natural language to robot behavior. Our observation is that these approaches can be situated on a spectrum ranging between two high-level approaches: mapping between language and a manually defined formal representation and mapping between language and high-dimensional vector spaces that translate directly to low-level robot policy. We define the advantages and limitations of each approach.
@@
Using a formal representation constrains the search space during training and inference and may require less training data. It also provides a natural framework for strong interpretability and formal safety guarantees. Well-defined model-checking tools exist to check whether the model of a system meets a given logical specification (Baier and Katoen, 2008). Formal methods can also synthesize correct-by-construction robot controllers given a logical specification and the system model and provide counterexamples to explain failure cases (Kress-Gazit et al., 2018). However, a formal representation constrains the space of possible models that can be learned, limiting the system’s ability to represent meanings for some things a person might say. With the advent of large language models, mapping between human language and a formal language is much easier; the research questions then need to focus on what formal language to use, where it comes from, and how it connects to the physical world. There are opportunities to more easily use existing representations such as the planning domain definition language (PDDL), linear temporal logic (LTL), or motion planners without collecting large training sets. Many approaches at the border, such as SayCan (Ichter et al., 2022), map language to a manually specified vocabulary of robot skills but give their formal language relatively little attention, despite it playing a critical role in the system.
@@
On the other hand, end-to-end approaches may require more data to train but are more flexible in representing a user’s intended meaning and translating it to robot behavior. Ng and Jordan (2001) observed that more structured generative models perform better with less training data. In contrast, discriminative models with more parameters can perform better when given lots of data because they make fewer assumptions about the structure of the learned model. Similarly, end-to-end neural embedding approaches require more training data but can generalize better than formal methods because they place fewer constraints on the learned model. True “pixels to torque” approaches (Wahlström et al., 2015) learn to produce motor torques directly from sensor input; however, many end-to-end approaches use some intermediate outputs such as end-effector poses or joint states. The challenge with end-to-end approaches is acquiring enough training data, processing this data, generalizing outside the training set, chaining policies together to produce long-term behaviors, explaining robot behaviors, and providing safety guarantees.
@@
We situate the robotic language grounding works on a spectrum in Figure 1 and describe more formal work in Section 2 and more end-to-end work in Section 3. More formal work uses discrete and structured representations that introduce bias into the learning. We begin with the most structured, abstract representations and successively review less structured, lower-level, continuous representations. Many of the implementations surveyed combine aspects of formal and end-to-end representations. In Section 4, we survey output representations for both approaches, review available datasets, and discuss the scope of natural language commands understood in theory and practice by different approaches. We also identify important open problems and recommend future research directions.
This new review paper focuses on the transformative role LLMs have played in this space. Tellex et al. (2020) is a review that surveys robot and language grounding work predating LLMs’ emergence. Zhang et al. (2023) reviews large language models more broadly used in human-robot interaction, including question answering, social robotics, and instruction following. Zeng et al. (2023b) reviews LLMs applied to robotics broadly, including related technologies, but does not focus on the spectrum from formal methods to high-dimensional vectors as this paper does. Wang et al. (2024) reviews applications of LLMs to robotics but does not situate the work on a spectrum and focuses on a broader set of tasks than command understanding. Our work, in contrast, focuses specifically on the problem of command understanding, situating work along a spectrum based on formal methods.
Table 1: Method Comparison
@@
LEFT (Hsu et al., 2023) Formal N/A FOL formula Low-level controllers Pick-and-Place (Real)
Lang2LTL (Liu et al., 2023b) Formal N/A LTL formula LTL-MDP planner, Low-level controllers Navigation (Real)
(Xie et al., 2023) Formal N/A PDDL goal Symbolic planner, Low-level controllers Navigation, Mobile manipulation (Sim)
LLM+P (Liu et al., 2023a) Formal N/A PDDL Symbolic planner, Low-level controllers Manipulation (Real)
SayCan (Ichter et al., 2022) Formal N/A predefined skill Low-level controllers Mobile manipulation (Real)
Code as Policies (Liang et al., 2023) Formal N/A Python code Python interpreter, Low-level controllers Pick-and-place, Mobile manipulation
@@
ProgPrompt (Singh et al., 2023) Formal N/A Python code Python interpreter, Low-level controllers Mobile manipulation (Sim)
@@
Voyager (Wang et al., 2023) Formal N/A Javascript Code Javascript interpreter, Low-level controllers Minecraft survival tasks (Sim)
ITP (Li et al., 2023a) Intermediate High-level action Python code Python interpreter, Low-level controllers Manipulation (Real)
@@
UniSim (Yang et al., 2023) Intermediate Image subgoals predefined skill Low-level controllers Manipulation (Real)
SuSIE (Black et al., 2023) Intermediate Image subgoals End-effector pose Low-level controllers Manipulation (Real)
VLP (Du et al., 2024) Intermediate Image subgoals End-effector pose Low-level controllers Manipulation (Real)
PaLM-E (Driess et al., 2023) End-to-End N/A predefined skill Low-level controllers Pick-and-place, Manipulation (Real)
VIMA (Jiang et al., 2023) End-to-End N/A predefined skill Low-level controllers Manipulation (Sim)
PerAct (Shridhar et al., 2023) End-to-End N/A End-effector pose Low-level controllers Manipulation (Real)
RT-1 (Brohan et al., 2023) End-to-End N/A End-effector pose Low-level controllers Manipulation (Sim, Real)
@@
RT-2 (Zitkovich et al., 2023) End-to-End N/A End-effector pose Low-level controllers Manipulation (Real)
RT-X (Padalkar et al., 2023) End-to-End N/A End-effector pose Low-level controllers Manipulation (Real)
@@ 2 Mapping from Natural Language to a Formal Representation
More structured methods tend to map more directly to certain natural language commands and express the goal or constraint directly rather than imperatively how to achieve it. Goal-based representations specify what state the world should be in but not what actions the robot should take to attain that state. In contrast, action-based representations specify a sequence of actions but not necessarily the goal or results of those actions. For example, a goal-based representation for “avoid the red room” might be a logical formula such as $\neg red\_room$ while an action-based expression might be $North;North;West;West;North;North$ (depending on the specific geometry of the environment). We order our review from high-level, abstract representations to low-level, concrete, fine-grained representations.
@@ 2.1 Logics
To train their language grounding system on diverse natural language commands, Pan et al. (2023) used LLMs to paraphrase structured English commands generated from algorithmically produced LTL formulas. Wang et al. (2021) and Patel et al. (2020) trained a semantic parser to map language commands to LTL formulas using weak supervision of execution trajectories without any LTL annotations. Lang2LTL (Liu et al., 2023b) is a modular system that uses LLMs to ground navigation commands to linear temporal logic (LTL) formulas and their propositions to physical landmarks in a given semantic map. The same system solved navigational tasks in indoor and outdoor environments without retraining on language data by harnessing pre-trained LLMs. Similarly, Hsiung et al. (2022) first translated commands to lifted LTL formulas and then grounded them to specific domains for better generalization. Other approaches (Fuggitti and Chakraborti, 2023; Chen et al., 2023) also used LLMs to translate natural language commands to logical representations but did not ground the formulas to a robot domain.
AutoTAMP (Chen et al., 2024) uses LLMs to translate task and state descriptions to signal temporal logic (STL) formulas (Maler and Nickovic, 2004) and correct syntax errors if detected. It then uses an STL planner to generate trajectories. By using an intermediate formal task specification, AutoTAMP outperforms LLM planners on tasks with geometric and temporal constraints in 2D domains.
Recent work also leveraged LLMs for grounding natural language commands to first-order logic (FOL) formulas. LEFT (Hsu et al., 2023) used an LLM to translate natural language queries to FOL programs, which a differentiable FOL executor executed. At the same time, a domain-specific grounding model grounded the symbols of the FOL program to various input modalities, e.g., 2D images and point clouds. The advantage of using first-order logic over LTL is capturing commands that use quantifiers and predicates for generalization. On the other hand, LTL provides a natural way to concisely represent temporally extended commands. Different logic representations can be formally composed to form a new, more expressive logic representation. These representations can enable direct mapping between abstract concepts in a language such as “avoid” and a precise, formal logical expression that guarantees task satisfaction.
@@ 2.2 Planning Domain Definition Language (PDDL)
The Planning Domain Definition Language (PDDL) (McDermott et al., 1998; Fox and Long, 2003; Edelkamp and Hoffmann, 2004; Kovacs, 2011) is a structured representation that defines a planning problem. It consists of a domain, which defines objects, predicates, and actions that govern the world’s rules, and a problem, a grounded problem instance with an initial state and a desired goal state. A symbolic planner takes as input the PDDL domain and problem and then outputs a plan (Helmert, 2006), i.e., a sequence of actions, to reach the goal state from the initial state. In this sense, it is a goal-based representation, but because it encodes actions and their effects, it can also be used imperatively. Recent works used LLMs to translate natural language descriptions of the world and the task to a PDDL representation of the planning problem, which a symbolic planner can then use. Compared to logical representations, PDDL provides a language of predicate states and transitions that allow goals to be translated into lower-level actions and skills but requires a full domain specification. People have defined PDDL domains for a wide variety of real-world applications. Konidaris et al. (2018) showed how a robot can learn symbols for low-level skills that are both necessary and sufficient to enable planning in PDDL.
@@
Guan et al. (2023) and Liu et al. (2023a) prompted an LLM to translate a natural language description of a problem into a complete PDDL problem definition, which was then fed into a symbolic planner together with a PDDL domain definition to find an optimal plan. These approaches outperform LLM planners and provide strong correctness guarantees from using a symbolic planner. Recent work has also created benchmarks for evaluating LLM’s ability to make plans concerning PDDL and other AI baselines (Valmeekam et al., 2023b, a), finding that LLMs achieve low success rates across domains in isolation. Still, they can improve the search process for underlying sound planners, leading to higher performance.
Unlike other works that use LLMs for AI planning, Silver et al. (2024) developed a generalized planner to synthesize Python programs to solve novel tasks by prompting LLMs with a domain specification and a few training tasks with an emphasis on satisfaction and efficiency. Their system also automatically detects planning errors and then re-prompts the LLM with feedback.
@@ 2.3 Code
Code as Policies (Liang et al., 2023) prompted an LLM with import statements, example code, and code comments that describe the desired policy to generate Python code executable directly on the robot. It solved manipulation and mobile manipulation tasks with API calls to Python libraries and predefined perception and control modules. ProgPrompt (Singh et al., 2023) applied a similar approach with additional assertion statements to recover from errors when reliable state tracking is available. Varley et al. (2024) developed a modular bi-arm system that employs an LLM to map natural language instructions to high-level API calls to manipulation skills powered by VLMs and a control module to solve three tabletop bimanual manipulation tasks. Their experiment results highlighted the benefits of modularity for ensuring safety, interpreting failures, and identifying modules to improve. Socratic Models (Zeng et al., 2023a) also demonstrated the code generation capabilities of LLMs for simulated pick-and-place tasks, provided with pre-trained perception modules and robot policies. ITP (Li et al., 2023a) used LLMs to generate high-level action sequences and then translated the action descriptions to predefined function calls using APIs for an object detector and robot policies to solve tabletop manipulation tasks. By storing completed high-level actions, ITP can replan given new commands at any step during the execution. Voyager (Wang et al., 2023) extended the code generation capabilities of LLMs to build a lifelong learning agent in Minecraft that continuously explores and learns a skill library of executable code.

### v2

@@ Abstract
With large language models, robots can understand language more flexibly and more capable than ever before. This survey reviews and situates recent literature into a spectrum with two poles: 1) mapping between language and some manually defined formal representation of meaning, and 2) mapping between language and high-dimensional vector spaces that translate directly to low-level robot policy. Using a formal representation allows the meaning of the language to be precisely represented, limits the size of the learning problem, and leads to a framework for interpretability and formal safety guarantees. Methods that embed language and perceptual data into high-dimensional spaces avoid this manually specified symbolic structure and thus have the potential to be more general when fed enough data but require more data and computing to train. We discuss the benefits and tradeoffs of each approach and finish by providing directions for future work that achieves the best of both worlds.
@@ 1 Introduction
Large language models (LLMs) have fueled a surge of interest in the problem of making robots understand natural language commands. Solving this problem requires mapping between words in language and actions or behaviors taken by the robot. Harnad (1990, 2007) defined the symbol grounding problem as constructing a mapping from symbols of a symbolic system or words in an utterance to sensorimotor substrates in the physical world. Large language models (LLMs) semantically represent concepts without explicit higher-order symbols beyond the words in the text, leading many to try end-to-end approaches for robotic language understanding. Yet many recent works in robotic language understanding leverage large language models hand-in-hand with formal symbolic representations. For example, Code as Policies Liang et al. (2022) generates Python code with predefined Python APIs. SayCan Ahn et al. (2022) grounds natural language commands to predefined discrete skills implemented using a deep neural network. Other approaches take a more end-to-end approach, such as VIMA Jiang et al. (2023), which learns a mapping from vision and language instructions to low-level robot actions such as joint states.
@@
This survey paper evaluates work that grounds natural language to robot behavior. We observe that these approaches can be situated on a spectrum ranging between two high-level approaches: mapping between language and a manually defined formal representation and mapping between language and high-dimensional vector spaces that translate directly to low-level robot policy. We define the advantages and limitations of each approach.
@@
Using a formal representation constrains the search space during training and inference and may require less training data. It also provides a natural framework for strong interpretability and formal safety guarantees. Well-defined model-checking tools exist to check whether the model of a system meets a given logical specification Baier and Katoen (2008). Formal methods can also synthesize correct-by-construction robot controllers given a logical specification and the system model and provide counterexamples to explain failure cases Kress-Gazit et al. (2018). However, a formal representation constrains the space of possible models that can be learned, limiting the system’s ability to represent the meanings of what a person may say. With the advent of large language models, mapping between human language and a formal language is much easier; the research questions then need to focus on what formal language to use, where it comes from, and how it connects to the physical world. There are opportunities to more easily use existing representations such as the planning domain definition language (PDDL), linear temporal logic (LTL), or motion planners without collecting large training sets. Many approaches at the border, such as SayCan Ahn et al. (2022), map language to a manually specified vocabulary of robot skills but give their formal language relatively little attention despite it playing a critical role in the system.
Figure 1: Approaches to representing natural language for robotics fall along a spectrum from more symbol-like representations to more continuous embedding-like representations. However, most approaches use a mixture of both. SayCan uses a fixed ontology of predefined skills but implements these as neural value functions conditioned on language.
@@
End-to-end approaches may require more data to train but are more flexible in representing a user’s intended meaning and translating it to robot behavior. Ng and Jordan (2001) observed that structured generative models perform better with less training data. In contrast, discriminative models with more parameters can perform better when given lots of data because they make fewer assumptions about the structure of the learned model. Similarly, end-to-end neural embedding approaches require more data but can generalize better than formal methods because they place fewer constraints on the learned model. True “pixels to torque” approaches Wahlström et al. (2015) learn to produce motor torques directly from sensor input; however, many end-to-end approaches use intermediate outputs such as end-effector poses or joint states. The challenge with end-to-end approaches is acquiring enough training data, processing this data, generalizing outside the training set, chaining policies together to produce long-term behaviors, explaining robot behaviors, and providing safety guarantees.
@@
We situate the robotic language grounding works on a spectrum in Figure 1 and describe more formal work in Section 2 and more end-to-end work in Section 3. More formal work uses discrete and structured representations that introduce bias into the learning. We begin with the most structured, abstract representations and successively review less structured, lower-level, continuous representations. Many of the implementations surveyed combine aspects of formal and end-to-end representations. In Section 4, we survey output representations for both approaches, review available datasets, and discuss the scope of natural language commands understood in theory and practice by different approaches. We conclude by identifying key open problems and recommending future research directions.
This new review paper focuses on the transformative role LLMs have played in this space. Other relevant reviews are:
  • Tellex et al. (2020) is a review that surveys robot and language grounding work predating LLMs’ emergence.
  • Zhang et al. (2023) reviews large language models more broadly used in human-robot interaction, including question answering, social robotics, and instruction following.
  • Zeng et al. (2023b) reviews LLMs applied to robotics broadly, including related technologies, but does not focus on the spectrum from formal methods to high-dimensional vectors as this paper does.
  • Wang et al. (2024) reviews applications of LLMs to robotics but does not situate the work on a spectrum and focuses on a broader set of tasks than command understanding.
Our work, in contrast, focuses specifically on the problem of command understanding, situating work along a spectrum based on formal methods.
@@
LEFT Hsu et al. (2023) Formal N/A FOL formula Low-level controllers Pick-and-Place (Real)
Lang2LTL Liu et al. (2023b) Formal N/A LTL formula LTL-MDP planner, Low-level controllers Navigation (Real)
Xie et al. (2023) Formal N/A PDDL goal Symbolic planner, Low-level controllers Navigation,
Mobile manipulation (Sim)
LLM+P Liu et al. (2023a) Formal N/A PDDL Symbolic planner, Low-level controllers Manipulation (Real)
SayCan Ahn et al. (2022) Formal N/A predefined skill Low-level controllers Mobile manipulation (Real)
Code as Policies Liang et al. (2022) Formal N/A Python code Python interpreter, Low-level controllers Pick-and-place, Mobile manipulation
@@
ProgPrompt Singh et al. (2023) Formal N/A Python code Python interpreter, Low-level controllers Mobile manipulation (Sim)
@@
Voyager Wang et al. (2023) Formal N/A Javascript Code Javascript interpreter, Low-level controllers Minecraft survival tasks (Sim)
ITP Li et al. (2023a) Intermediate High-level action Python code Python interpreter, Low-level controllers Manipulation (Real)
@@
UniSim Yang et al. (2023) Intermediate Image subgoals predefined skill Low-level controllers Manipulation (Real)
SuSIE Black et al. (2023) Intermediate Image subgoals End-effector pose Low-level controllers Manipulation (Real)
VLP Du et al. (2023) Intermediate Image subgoals End-effector pose Low-level controllers Manipulation (Real)
PaLM-E Driess et al. (2023) End-to-End N/A predefined skill Low-level controllers Pick-and-place, Manipulation (Real)
VIMA Jiang et al. (2023) End-to-End N/A predefined skill Low-level controllers Manipulation (Sim)
PerAct Shridhar et al. (2023) End-to-End N/A End-effector pose Low-level controllers Manipulation (Real)
RT-1 Brohan et al. (2023a) End-to-End N/A End-effector pose Low-level controllers Manipulation (Sim, Real)
@@
RT-2 Brohan et al. (2023b) End-to-End N/A End-effector pose Low-level controllers Manipulation (Real)
RT-X O’Neill et al. (2024) End-to-End N/A End-effector pose Low-level controllers Manipulation (Real)
Table 1: Method Comparison
@@ 2 Mapping from Natural Language to a Formal Representation
More structured methods often map directly to certain natural language commands and express the goal or constraint directly rather than imperatively how to achieve it. Goal-based representations specify what state the world should be in but not what actions the robot should take to attain that state. In contrast, action-based representations specify a sequence of actions but not necessarily the goal or results of those actions. A goal-based representation for “avoid the red room” might be a logical formula such as $\neg red\_room$ while an action-based expression might be $North;North;West;West;North;North$ (depending on the specific geometry of the environment). We order our review from high-level, abstract representations to low-level, concrete, fine-grained representations.
@@ 2.1 Logics
To train their language grounding system on diverse natural language commands, Pan et al. (2023) used LLMs to paraphrase structured English commands generated from algorithmically produced LTL formulas. Wang et al. (2021) and Patel et al. (2020) trained a semantic parser to map language commands to LTL formulas using weak supervision of execution trajectories without any LTL annotations. Lang2LTL Liu et al. (2023b) is a modular system that uses LLMs to ground navigation commands to linear temporal logic (LTL) formulas and their propositions to physical landmarks in a given semantic map. The same system solved navigational tasks in indoor and outdoor environments without retraining on language data by harnessing pre-trained LLMs. Similarly, Hsiung et al. (2022) first translated commands to lifted LTL formulas and then grounded them to specific domains for better generalization. Other approaches Fuggitti and Chakraborti (2023); Chen et al. (2023) also used LLMs to translate natural language commands to logical representations but did not ground the formulas to a robot domain.
AutoTAMP Chen et al. (2024) uses LLMs to translate task and state descriptions to signal temporal logic (STL) formulas Maler and Nickovic (2004) and correct syntax errors if detected. It then uses an STL planner to generate trajectories. By using an intermediate formal task specification, AutoTAMP outperforms LLM planners on tasks with geometric and temporal constraints in 2D domains.
Recent work also leveraged LLMs for grounding natural language commands to first-order logic (FOL) formulas. LEFT Hsu et al. (2023) used an LLM to translate natural language queries to FOL programs, which a differentiable FOL executor executed. At the same time, a domain-specific grounding model grounded the symbols of the FOL program to various input modalities, e.g., 2D images and point clouds. The advantage of using first-order logic over LTL is capturing commands that use quantifiers and predicates for generalization. On the other hand, LTL provides a natural way to concisely represent temporally extended commands. Different logic representations can be formally composed to form a new, more expressive logic representation. These representations can enable direct mapping between abstract concepts in a language such as “avoid” and a precise, formal logical expression that guarantees task satisfaction.
@@ 2.2 Planning Domain Definition Language (PDDL)
The Planning Domain Definition Language (PDDL) McDermott et al. (1998); Fox and Long (2003); Edelkamp and Hoffmann (2004); Kovacs (2011) is a structured representation that defines a planning problem. It consists of a domain, which defines objects, predicates, and actions that govern the world’s rules, and a problem, a grounded problem instance with an initial state and a desired goal state. A symbolic planner takes as input the PDDL domain and problem and then outputs a plan Helmert (2006), i.e., a sequence of actions, to reach the goal state from the initial state. In this sense, it is a goal-based representation, but because it encodes actions and their effects, it can also be used imperatively. Recent works used LLMs to translate natural language descriptions of the world and the task to a PDDL representation of the planning problem, which a symbolic planner can then use. Compared to logical representations, PDDL provides a language of predicate states and transitions that allow goals to be translated into lower-level actions and skills but requires a full domain specification. People have defined PDDL domains for a wide variety of real-world applications. Konidaris et al. (2018) showed how a robot can learn symbols for low-level skills that are both necessary and sufficient to enable planning in PDDL.
@@
Guan et al. (2023) and Liu et al. (2023a) prompted an LLM to translate a natural language description of a problem into a complete PDDL problem definition, which was then fed into a symbolic planner together with a PDDL domain definition to find an optimal plan. These approaches outperform LLM planners and provide strong correctness guarantees from using a symbolic planner. Recent work has also created benchmarks for evaluating LLM’s ability to make plans concerning PDDL and other AI baselines Valmeekam et al. (2023a, b), finding that LLMs achieve low success rates across domains in isolation. Still, they can improve the search process for underlying sound planners, leading to higher performance.
Unlike other works that use LLMs for AI planning, Silver et al. (2024) developed a generalized planner to synthesize Python programs to solve novel tasks by prompting LLMs with a domain specification and a few training tasks emphasizing satisfaction and efficiency. Their system also automatically detects planning errors and then re-prompts the LLM with feedback.
@@ 2.3 Code
Code as Policies Liang et al. (2022) prompted an LLM with import statements, example code, and code comments that describe the desired policy to generate Python code executable directly on the robot. It solved manipulation and mobile manipulation tasks with API calls to Python libraries and predefined perception and control modules. ProgPrompt Singh et al. (2023) applied a similar approach with additional assertion statements to recover from errors when reliable state tracking is available. Varley et al. (2024) developed a modular bi-arm system that employs an LLM to map natural language instructions to high-level API calls to manipulation skills powered by VLMs and a control module to solve three tabletop bimanual manipulation tasks. Their experiment results highlighted the benefits of modularity for ensuring safety, interpreting failures, and identifying modules to improve. Socratic Models Zeng et al. (2023a) also demonstrated the code generation capabilities of LLMs for simulated pick-and-place tasks, provided with pre-trained perception modules and robot policies. ITP Li et al. (2023a) used LLMs to generate high-level action sequences and then translated the action descriptions to predefined function calls using APIs for an object detector and robot policies to solve tabletop manipulation tasks. By storing completed high-level actions, ITP can replan given new commands at any step during the execution. Voyager Wang et al. (2023) extended the code generation capabilities of LLMs to build a lifelong learning agent in Minecraft that continuously explores and learns a skill library of executable code.